Trace the curve y^2 = (x +1) (x −1)^2 by showing all the properties you use to trace it.
1
Expert's answer
2019-08-19T08:23:39-0400

Explanation

Since the equation does not change when we replace y with –y, then the curve has reflective symmetry with x-axis.

We can find intersection points of the curve with x-axis from equation y = 0. Hence, (x+1)(x−1)2 = 0, or x = -1, x = 1. From equation x = 0 we find intersection points of the curve with y-axis: y2 = 1, or y = -1, y = 1.

Because x+1 = y2/(x-1)2 (x≠1), then x+1≥0, x≥ -1.

Let us rewrite the equation

y2-(x+1) (x−1)2 = 0, that is (y-(x-1)"\sqrt{x+1}" )(y+(x-1)"\sqrt{x+1}" ) = 0.

Hence the curve has two branches y=(x-1)"\sqrt{x+1}", y=-(x-1)"\sqrt{x+1}" .

Because y/x=(1-1/x)"\sqrt{x+1}" →"\infin" as x tends "\infty", these branches have not asymptotes.

A branch y=(x-1)"\sqrt{x+1}" is the graph of the function y=(x-1)"\sqrt{x+1}". For analyzing this function we find its derivative. It is equal to y'="(3x+1)\over2\sqrt{x+1}". The derivative is negative for x<-1/3, it is positive for x>-1/3 and equals to zero at x=-1/3. Hence the function is decreasing over the interval ]-1, -1/3[ and is increasing over the interval ]-1/3,"\infin" [. At the point x=-1/3 it has the minimal value ymin = "-4\sqrt{2}\over3\sqrt{3}".

Since the second derivative of the function y'' = "\sqrt{x+1}(3x+5)\over(2x+2)^2" is positive over the interval ]-1,"\infin"[, then the graph of function is concave up on the interval ]-1,"\infin" [.

Using the mentioned properties we can trace the curve.